### v2

# Partition function zeros of quantum many-body systems

Abstract
We present a method for calculating the Yang-Lee partition function zeros of a translationally invariant model of lattice fermions, exemplified by the Hubbard model. The method rests on a theorem involving the single electron self-energy $\Sigma_{\sigma}(\vec{k},i\omega_{n})$ in the imaginary time Matsubara formulation. The theorem maps the Yang-Lee zeros to a set of wavevector and spin labeled virtual energies $\xi_{\vec{k}{\sigma}}$. These, thermodynamically derived virtual energies, are solutions of equations involving the self-energy at corresponding $\vec{k}{\sigma}$’s. Examples of the method in simplified situations are provided.
1 Introduction
C. N. Yang and T. D. Lee launched an important and fruitful direction in condensed matter physics, by highlighting the significance of the zeros of the grand partition function of many-body systems[1]. Lee and Yang[2], soon thereafter, showed that for Ising ferromagnets, the zeros can be located exactly and lie on the unit circle in the complex fugacity plane. They showed that the density of the zeros closest to the real line provides fundamental insight into the critical singularities near a phase transition in 2 and higher dimensions. The work of T. Asano[3] gave an elegant reformulation of the Lee-Yang results using the method of “contractions”, and also enabled the surmounting of the added difficulties due to quantum mechanics in the ferromagnets[4]. The methodology used for these results was streamlined by D. Ruelle and F. Dyson[5], leading to several important further results over the intervening years[6, 7], reviewed in [8]. The currently existing methods unfortunately do not appear to be generalizable for antiferromagnets and for broader classes of systems, including the ones considered here.
Our interest is in the study of quantum many-body systems, such as the Hubbard model. Locating the zeros and extracting meaningful information for these systems has proceeded at a sedate pace. Here the major difficulty is the absence of systematic analytical methods. For numerical studies, a combination of the very high precision required for locating the zeros and the exponential growth of the Hilbert spaces with the system size, hamper direct enumeration techniques. Quantum Monte Carlo (QMC) methods for studying the repulsive Hubbard model with a complex chemical potential have been used earlier [9, 10, 11, 12]. However the full spectrum of zeros seems difficult to obtain from QMC. There have also been recent studies of partition function zeros for related quantum many-body models in a variety of contexts, these include decoherence studies[13], quantum quenches[14] and in quantum computation [15]. These indicate a renewal of interest in this important subject.
The main objective of this paper is to introduce a new analytical method for locating the zeros of the grand partition function of the Hubbard model (see Eq. (5))- in any dimension. This method requires a knowledge of the self energy in the Matsubara imaginary time framework[16]. Perturbative knowledge of the self-energy may be used to give approximate location of the zeros. The method generalizes to nearby models in a straightforward way. The a few simple examples are provided below.
1.1 Heuristic perturbative approach to Yang-Lee zeros
Let us consider the Hubbard model Eq. (5) as an illustrative example. A perturbative approach is to expand the (grand) partition function as a series in the coupling constant [17] $Z={\text{Tr}}e^{-\beta({\cal H}_{0}+V)}=Z_{0}(1-\beta\langle T_{\tau}V\rangle_{0})+{\cal O}(U^{2})$, where $Z_{0}=\prod_{\alpha}(1+ze^{-\beta\varepsilon_{\alpha}})^{2}$ is the non-interacting result with the fugacity $z=e^{\beta\mu}$, $T_{\tau}$ the imaginary-time ordering operator, $\beta=\frac{1}{k_{B}T}$ and the correction term due to interactions is found by taking the non-interacting thermal average of the interaction term, so that
$Z(z)=\prod_{\alpha}(1+ze^{-\beta\varepsilon_{\alpha}})^{2}\times\left[1-\frac{\beta U}{N_{s}}\left(\sum_{\alpha}\frac{z}{z+e^{\beta\varepsilon_{\alpha}}}\right)^{2}+{\cal O}(U^{2})\right],$ (1)
where the spin degeneracy leads to the square in the second term. For simplicity let us assume that all the energies $\varepsilon_{\alpha}$ are distinct. A simple minded estimate of the location of the zeros can be made as follows. Since the zeros of $Z_{0}$ are located at $z_{\alpha}\to-e^{\beta\varepsilon_{\alpha}}$, we might expect that the zeros of $Z$ are continuously connected to these and hence try the expression $z_{\alpha}=-e^{\beta(\varepsilon_{\alpha}+\Delta_{\alpha})}$ where the energy shift $\Delta_{\alpha}$ vanishes as $U\to 0$. With a fixed $\alpha$ we substitute into Eq. (1) and the vanishing condition imposed on $Z$, thus requiring
$\sum_{\gamma}\left(\frac{1}{1-e^{\beta(\varepsilon_{\gamma}-\varepsilon_{\alpha}-\Delta_{\alpha})}}\right)=\pm\sqrt{\frac{N_{s}}{\beta U}}.$ (2)
Solving to low orders we get
$\Delta_{\alpha}=\pm\sqrt{\frac{Uk_{B}T}{N_{s}}}+\frac{U}{2N_{s}}\sum_{\gamma\neq\alpha}\coth{\frac{\beta}{2}(\varepsilon_{\gamma}-\varepsilon_{\alpha})}+\ldots$ (3)
From this expression we infer that for $U>0$ the energy shifts are real, and the fugacities $z_{\alpha}$ are on the negative real axis. On the other hand for any $U<0$ the energy shifts are complex numbers occurring in complex conjugate pairs, and the fugacities $z_{\alpha}$ lie on a curve in the negative half plane which is symmetric about the real axis. In Fig. (1) we present the roots for the 4-site Hubbard model, where the partition function was calculated using the software package DiracQ[18] employing symbolic methods. We see that the exact roots behave in accordance with the above observation. The change in the nature of roots- from real for repulsive to non-real for attractive interactions is pronounced, and easily visible. This provides a graphic demonstration of Dyson’s comments[19] on the nature of perturbation theory in many-body systems, and is discussed further below.
Figure 1: Yang-Lee zeros of the partition function of the 4 site Hubbard model at T=1: TopLeft with U=-0.5 TopRight with U=-0.25 BottomLeft with U=+0.25 and BottomRight with U=0.5. The partition function was calculated using the symbolic calculation programme DiracQ[18]. From these figures we see that the attractive case leads to complex fugacities lying on a parabolaic curve- while the repulsive case leads to real fugacities. This is as expected from Eq. (3).
We therefore obtain some insight into the roots of the Hubbard model from the simple perturbation theory outlined above. However the theoretical argument is not completely satisfactory, since it mixes in factors of very different orders in $N_{s}$ in the expression Eq. (1), or its higher order generalizations. It is highly desirable to have a formulation where the fugacity-which is of the ${\cal O}(1)$ in $N_{s}$- is found from equations where all terms are of the same, i.e. ${\cal O}(1)$ order. We provide such a framework in this work.
1.2 Plan of the paper
In Section(2) we provide basic definitions of the Hamiltonian, the grand partition function and the fugacity. We define the virtual energies $\xi_{\vec{k}{\sigma}}$, and note some properties of these obtained from general considerations. In Section(3) we define the Matsubara finite temperature Greens functions, their un-normalized versions and the self-energy. In Section(4) the main theorem of this work is proposed and established. The equation determining the virtual energy in terms of the self-energy is recorded in Eq. (28). In Section(5) we illustrate the application of the theorem to the problem of an exactly solvable single site Hubbard model. Here the role of lifting of degeneracy of the single site energy is also studied in detail. The exact self energy is expanded to ${\cal O}(U^{2})$, and to this order the perturbative results for $\xi$ are shown to agree with the exact ones. This somewhat long section is mainly addressed to readers interested in details, and may be skipped by others. In Section(6) we use the first order (Hartree) approximation to the self energy and extract the virtual energies $\xi_{\vec{k}{\sigma}}$ to ${\cal O}(U)$ for both signs of the interaction for the 1-d Hubbard ring of length 6. This is familiar as the idealized Benzene molecule. These virtual zeros are compared with exact numerical values obtained by computing the grand partition function exactly using symbolic methods. In Section(7) we note a few conclusions and prospects for further work.
In Appendix(A) we recall the well known results of Luttinger and Ward[20] for the grand partition function. Their result expresses the result of an all-orders summation of perturbation theory. We observe that the key function $\Xi(\mu)$ in Eq. (27)- whose zeros are all the virtual energies- is a multiplicative factor of the Luttinger-Ward partition function.
2 Basic Hamiltonian and the partition function
Let us consider an interacting fermi system described by the (grand-canonical) Hamiltonian
${\cal H}$ $=$ $H-\mu\,{\cal N}={\cal H}_{0}+V$
${\cal H}_{0}$ $=$ $\sum_{\vec{k}_{\alpha}{\sigma}}\varepsilon_{k_{\alpha},{\sigma}}\,C^{\dagger}_{\vec{k}_{\alpha}{\sigma}}C_{\vec{k}_{\alpha}{\sigma}}-\mu{\cal N},$ (4)
where the (conserved) fermion number operator ${\cal N}=\sum_{k,{\sigma}}C^{\dagger}_{k{\sigma}}C_{k{\sigma}}$, $\mu$ the chemical potential, and the Hubbard Hamiltonian
${H}=\sum_{\vec{k}_{\alpha}{\sigma}}\varepsilon_{k_{\alpha},{\sigma}}\,C^{\dagger}_{\vec{k}_{\alpha}{\sigma}}C_{\vec{k}_{\alpha}{\sigma}}+U\sum_{i}n_{i\uparrow}n_{i\downarrow}.$ (5)
Here $\vec{k}_{\alpha}$ are the $L^{d}$ box-quantized wave vectors. The kinetic energy term is the Fourier transform of the hopping Hamiltonian $T=-\sum_{\vec{r},{\sigma},\vec{\eta}}(t_{\vec{\eta},{\sigma}}C^{\dagger}_{\vec{r}{\sigma}}C_{\vec{r}+\vec{\eta}{\sigma}}+h.c.)$, and $\vec{\eta}$ is the nearest neighbour vector on a d-dimensional hypercubic lattice. In the usual case the hopping matrix element $t_{\vec{\eta},{\sigma}}$ is independent of the spin ${\sigma}$. The generalized ${\sigma}$ dependent case is useful for studying the effect of lifting the spin-degeneracy on the roots as discussed below. More general models can be obtained by replacing the interaction term with an appropriate $V$ and the energy dispersion chosen suitably. We have assumed that the system has translation invariance- and study a finite lattice (with lattice constant unity) with $N_{s}=L^{d}$ sites.
For the finite system under consideration the total number of (spin half) particles varies in the range $0\leq N\leq 2N_{s}$. With the Boltzman constant $k_{B}=1$ and $\beta=\frac{1}{T}$ we trace over all particle numbers and define the grand partition function:
$Z(\mu)={\text{Tr}}\,e^{-\beta(H-\mu{\cal N})}=\sum_{n,N}e^{-\beta E_{n}+\beta\mu N_{n}}=\sum_{N=0}^{2N_{s}}z^{N}Z_{N}(\beta),$ (6)
where $E_{n}$ and $N_{n}$ are the eigenvalues of $H$ and particle number for the state $n$ (detailed definitions are given after Eq. (29)), $Z_{N}$ is the canonical partition function for $N$ particles and $z$ is the fugacity
$z(\mu)=e^{\beta\mu}.$ (7)
$Z(\mu)$ is thus a polynomial in the fugacity $z$ of degree $2N_{s}$, and by the fundamental theorem of algebra, it has exactly $2N_{s}$ zeros in the complex $z$ plane- and hence expressible as $Z(\mu)=c_{0}\prod_{\alpha=1}^{2N_{s}}(z-z_{\alpha})$, where $z_{\alpha}$ are the Yang-Lee zeros of the given model. We will find it more convenient to work in the complex $\mu$ plane rather than with the fugacity $z$. We note the periodicity $z(\mu)=z(\mu+i\frac{2\pi}{\beta})$ and therefore the complex $\mu$ plane can be decomposed into infinite equivalent strips of width $i\frac{2\pi}{\beta}$. We restrict our attention to a conveniently chosen strip
${\cal S}=\left\{-\frac{2\pi}{\beta}<\Im m\,\mu\leq 0,\;\;\Re e\,\mu\in[-\infty,\infty]\right\}.$ (8)
We remark further on the rationale underlying this convention below Eq. (11).
2.1 Definition of the virtual energies $\xi_{\vec{k}{\sigma}}$
We will write the partition function in the form
$Z(\mu)=\prod_{\vec{k}{\sigma}}(1+\frac{e^{\beta\mu}}{e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{U}{2})}}).$ (9)
With this convention- explained below- the roots of $Z$ occur at
$\mu$$\to$ $-i\frac{\pi}{\beta}+\frac{U}{2}+\xi_{\vec{k}{\sigma}},\mbox{ or equivalently }$
$z_{\alpha}$$\to$ $-e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{U}{2})}$ (10)
where Eq. (10) anticipates a mapping $\alpha\leftrightarrow\vec{k}{\sigma}$, between the unlabeled Yang-Lee zeros $z_{\alpha}$ and the labeled virtual energies $\xi$. We remark that while Eq. (9) has the same form as that of an ideal Fermi gas with energies $\xi_{\vec{k}{\sigma}}+\frac{U}{2}$, the latter are non-trivial complex functions of $T,U$. For an interacting system $\xi_{\vec{k}{\sigma}}$ are not viewed as eigenvalues of any Hermitean Hamiltonian, and are allowed to be non-real. For this reason these may be regarded as thermodynamically derived virtual energies.
The imaginary part of $\xi_{\vec{k}{\sigma}}$ is constrained by
$-\frac{\pi}{\beta}<\Im m\,\xi_{\vec{k}{\sigma}}\leq\frac{\pi}{\beta}.$ (11)
The convention used in defining the strip ${\cal S}$ in Eq. (8), and the range of the imaginary part of $\xi$ in Eq. (11), are such that complex conjugate roots in the fugacity variable $z$ Eq. (7) lead to complex conjugate energies $\xi$. The partition function being a polynomial in $z$ with real coefficients, the roots in $z$ are either real or come in complex conjugate pairs. It follows that the virtual energies are either real or they occur in complex conjugate pairs, i.e. if $\xi$ is a virtual energy, then so is $\xi^{*}$.
The virtual energies $\xi$’s are dependent on the temperature $T$, and also the interaction strength $U$. We also remark that if the virtual energies $\xi$ are real and their density of states is non-singular in a continuous interval of energies, then the system would behave like an almost ideal Fermi liquid for densities where $\mu$ lives in the same range. Such a situation seems to be realized in the repulsive Hubbard model at low $T$ studied here. It is tempting to view Landau’s Fermi liquids [21, 22, 17] as behaving in precisely such a fashion, with the energies $\xi_{\vec{k}{\sigma}}$ being closely related to Landau’s quasiparticle energies [21].
We next record some general properties of the virtual energies for the Hubbard model.
  • On bipartite lattices there is an invariance under a simultaneous particle hole transformation for both directions of the spin, it is expressible in real space as $C_{i{\sigma}}\leftrightarrow e^{i\pi\phi_{i}}C^{\dagger}_{i{\sigma}}$ with a suitable phase factor $\phi_{i}$. Collecting the extra coefficients generated by the transformation, we obtain an identity
    ${Z}(\mu)=e^{2\beta\mu N_{s}}e^{-\beta UN_{s}}{Z}(U-\mu).$ (12)
    This implies that the partition function zeros arise in particle-hole symmetric pairs $\{\mu,U-\mu\}$, or equivalently the virtual energies come in pairs, i.e.
    $\xi_{k{\sigma}}=-\xi_{k^{\prime}{\sigma}},$ (13)
    for suitable momentum partners $k$ and $k^{\prime}$.
  • In the fully occupied subspace (i.e. particle number $N=2N_{s}$ sector), there is only a single state. Here the kinetic energy is inoperative due to the “jammed” nature of the state, and hence it is simple to compute the corresponding canonical partition function $Z_{2N_{s}}=e^{-\beta UN_{s}}$. Comparing with the coefficient of $e^{\beta N_{s}\mu}$ in Eq. (9), we find the first sum-rule:
    $\sum_{\vec{k}{\sigma}}\xi_{\vec{k}{\sigma}}=0.$ (14)
    The imaginary parts of $\xi$, if present, cancel out by the complex conjugate property of the roots. It is readily seen that Eq. (14) is satisfied when the $\xi$’s satisfy the condition of particle hole symmetric pairs Eq. (13) by summing over all $\{\vec{k}{\sigma}\}$.
  • A second sumrule can be readily found by comparing the results of the single particle sector $N=1$, i.e the coefficient of $e^{\beta\mu}$ in Eq. (9). Its exact value is found using the observation that the two-particle interaction $U$ is inoperative in this sector. We find the second sum-rule:
    $\sum_{\vec{k}{\sigma}}e^{-\beta(\xi_{\vec{k}{\sigma}}+\frac{U}{2})}=\sum_{\vec{k}{\sigma}}e^{-\beta\varepsilon_{\vec{k}{\sigma}}}.$ (15)
  • Further sum-rules can be found by comparing other powers of $e^{\beta\mu}$, but they require exact knowledge of the canonical partition functions $Z_{2},Z_{3}\ldots$, and are therefore less tractable.
3 The Greens function and Self energy in Matsubara frequencies
It is useful to consider the unnormalized thermal Greens function, signified by an overbar
$\bar{G}_{\sigma}(\vec{k},\tau|\mu)=-\left({\text{Tr}}\,e^{-\beta{\cal H}}\left(T_{\tau}C_{\vec{k}{\sigma}}(\tau)C^{\dagger}_{\vec{k}{\sigma}}(0)\right)\right),$ (16)
where $-\beta<\tau\leq\beta$, and $T_{\tau}$ denotes the (Fermionic) imaginary-time ordering operator. For clarity we display the dependence on the chemical potential for $G,\Sigma,\ldots$, as in Eq. (16). For any operator ${\cal Q}$ the standard imaginary time Heisenberg picture time dependence is defined through ${\cal Q}(\tau)=e^{\tau{\cal H}}{\cal Q}e^{-\tau{\cal H}}$. This Greens function is related to the conventional normalized Greens function $G$ [16, 17] through
$\bar{G}_{\sigma}(\vec{k},\tau|\mu)=Z(\mu)G_{\sigma}(\vec{k},\tau|\mu).$ (17)
[…]
$(-\partial_{\tau}+\mu-\varepsilon_{\sigma}(k))\bar{G}_{\sigma}(k,\tau|\mu)=\delta(\tau)Z(\mu)+\bar{\cal A}_{\sigma}(\vec{k},\tau|\mu)$ (18)
where
$\bar{\cal A}(\vec{k},\tau|\mu)=-{\text{Tr}}\,e^{-\beta{\cal H}}\left(T_{\tau}a_{\vec{k}{\sigma}}(\tau)C^{\dagger}_{\vec{k}{\sigma}}(0)\right)$ (19)
with
$a_{\vec{k}{\sigma}}\equiv[C_{\vec{k}{\sigma}},V].$ (20)
Note that we can also write the un-normalized $\bar{\cal A}=Z(\mu){\cal A}$, in analogy with the corresponding G’s. Taking the Matsubara-Fourier series of both sides we get
$\left\{i\omega_{r}+\mu-\varepsilon_{\sigma}(k)\right\}\bar{G}_{\sigma}(\vec{k},i\omega_{r}|\mu)=Z(\mu)+\bar{\cal A}_{\sigma}(\vec{k},i\omega_{r}|\mu).$ (21)
Defining the Dyson self energy $\Sigma$ as
$\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)=\frac{\bar{\cal A}_{\sigma}(\vec{k},i\omega_{r}|\mu)}{\bar{G}_{\sigma}(\vec{k},i\omega_{r}|\mu)}$ (22)
this leads to the un-normalized version of Dyson’s equation
$\left\{i\omega_{r}+\mu-\varepsilon_{\sigma}(k)-\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)\right\}\,\bar{G}_{\sigma}(\vec{k},i\omega_{r}|\mu)=Z(\mu).$ (23)
It should be noted that in $\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)$, every occurrence of $i\omega_{r}$ is in the combination $i\omega_{r}+\mu$. In terms with $r\neq 0$ we note that $\omega_{r}-\omega_{0}=r\frac{2\pi}{\beta}$, this shift can be absorbed into $\mu$ and using the periodicity $Z(\mu)=Z(\mu+ir\frac{2\pi}{\beta})$. We also note that Eq. (302122) implies the periodicity in the complex $\mu$ plane for integer $s$
$\bar{G}_{\sigma}(\vec{k},i\omega_{r}|\mu+is\frac{2\pi}{\beta})=\bar{G}_{\sigma}(\vec{k},i\omega_{r+s}|\mu),$ (24)
and similar relations for $\Sigma_{\sigma}(\vec{k},i\omega_{r})$. Therefore studying $\bar{G}_{\sigma}(\vec{k},i\omega_{r}|\mu)$ with $r\neq 0$ does not lead to anything new and in the following we confine ourselves to $r=0$. Displaying the $\mu$ dependence of $G$ and $\Sigma$ explicitly and defining $\Psi$ through
$\Psi_{\vec{k}\,{\sigma}}(\mu)$$\equiv$ $G^{-1}_{{\sigma}}(\vec{k},i\frac{\pi}{\beta}|\mu)$ (25)
$=$ $\mu+i\frac{\pi}{\beta}-\varepsilon_{\sigma}({\vec{k}})-\Sigma_{\sigma}(\vec{k},i\frac{\pi}{\beta}|\mu),$
we write Eq. (23) as
$\Psi_{\vec{k}\,{\sigma}}(\mu)\times\bar{G}_{\sigma}(\vec{k},i\frac{\pi}{\beta}|\mu)=Z(\mu).$ (26)
The total number of such equations is $2N_{s}$, including the 2 spin projections and the $N_{s}$ values of $\vec{k}$.
4 Theorem for Locating the Yang-Lee zeroes of $Z(\mu)$.
We now discuss the connection between Yang-Lee zeros of $Z(\mu)$ and the vanishing of $\Psi_{\vec{k}{\sigma}}(\mu)$ (see Eq. (25)). Recalling Eq. (8) we will establish below
Theorem.
For a fixed $(\vec{k}\,{\sigma})$, if a complex number $\mu^{*}\in{\cal S}$ is a zero of $\Psi_{\vec{k}{\sigma}}(\mu)$ (i.e. $\Psi_{\vec{k}{\sigma}}(\mu^{*})=0$), then $\mu^{*}$ is also a zero of the partition function (i.e. $Z(\mu^{*})=0)$.
Remarks: We record a few relevant comments below.
  • In principle the self-energy $\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)$ can be calculated to arbitrary orders in the interaction, using the standard rules of the Feynman-Dyson-Matsubara perturbation theory in the imaginary time[16, 17, 20]. Given an approximation for $\Sigma$ to a certain order in the coupling, the theorem can be used to find the corresponding (unlabeled) Yang-Lee zero of the partition function, or more precisely, the momentum-labeled virtual energy $\xi_{\vec{k}{\sigma}}$.
  • If the non-interacting virtual energies (i.e. the band energies $\varepsilon_{\vec{k}{\sigma}}$) are completely non-degenerate, then using continuity in the coupling, all zeros of $Z$ can be located perturbatively from the theorem by varying $\vec{k},{\sigma}$.
  • Eq. (26) also admits the possibility that $\mu^{*}$, a root of $Z$, satisfies $\bar{G}_{\sigma}(\vec{k},i\frac{\pi}{\beta}|\mu^{*})=0$ for a fixed $\vec{k}{\sigma}$, while $\Psi_{\vec{k}{\sigma}}(\mu^{*})\neq 0$. This type of root is much less amenable to a systematic perturbative treatment since $\bar{G}$, unlike $\Psi_{\vec{k}{\sigma}}$, is generally a function of almost the entire set of $\vec{k}{\sigma}$. However we may bypass this class of roots- by studying the union over the $\vec{k},{\sigma}$ of the class of roots for each individual $\vec{k},{\sigma}$’s.
  • With the non-degeneracy assumption noted above, our result is that all zeros of $Z$ coincide with the zeros of a function $\Xi(\mu)$ defined by
    $\Xi(\mu)=\prod_{\vec{k}{\sigma}}\Psi_{\vec{k}{\sigma}}(\mu).$ (27)
  • When the single electron levels $\varepsilon_{\vec{k}{\sigma}}$ have degeneracies due to spin invariance or parity, this case can be viewed as a limiting case of the non-degenerate one. In the discussion following Eq. (68) we remark on how the degenerate case, treated to lowest order in $U$, allows for two perturbed solutions arising from a single (unperturbed) degenerate one, so that one can assign them one virtual energy to each spin projection.
  • Using Eq. (10) and Eq. (25) we summarize the final equation determining the virtual energy $\xi_{\vec{k}{\sigma}}$ as
    $\xi_{\vec{k}{\sigma}}=\varepsilon_{\vec{k}{\sigma}}-\frac{U}{2}+\Sigma_{{\sigma}}(\vec{k},i\frac{\pi}{\beta}|\xi_{k{\sigma}}+\frac{U}{2}-i\frac{\pi}{\beta}).$ (28)
Proof: To prove the theorem, a glance at Eq. (26) shows that it is sufficient to show that $\bar{G}$ is a holomorphic function of $\mu$ in the strip ${\cal S}$ (Eq. (8)) (i.e. without any poles or other singular points). We can establish this property using the standard eigenbasis representation
$\bar{G}_{\sigma}(\vec{k},i\omega_{r}|\mu)=\sum_{n,m}\frac{e^{-\beta\epsilon_{n}}+e^{-\beta\epsilon_{m}}}{\epsilon_{n}-\epsilon_{m}+i\frac{\pi}{\beta}(2r+1)}|\langle n|C_{\sigma}(\vec{k})|m\rangle|^{2}$ (29)
where $\left\{|n\rangle\right\}$ denotes a complete set of (Fock) states ${\cal H}|n\rangle=\epsilon_{n}|n\rangle,H|n\rangle=E_{n}|n\rangle,{\cal N}|n\rangle=N_{n}|n\rangle$, so that $\epsilon_{n}=E_{n}-\mu N_{n}$. This spectral representation follows from Eq. (16) by inserting complete sets of eigenstates followed by a Fourier transform from $\tau$ to the Matsubara frequencies $\omega_{r}$. Recall that the destruction operator $C_{\sigma}$ acting to the right either destroys the state or decreases the number of particles by 1, i.e. $N_{n}-N_{m}=-1$. Making the $\mu$ dependence explicit we write
$\frac{e^{-\beta\epsilon_{n}}+e^{-\beta\epsilon_{m}}}{\epsilon_{n}-\epsilon_{m}+i\frac{\pi}{\beta}(2r+1)}=e^{\beta(\mu N_{n}-E_{n})}\frac{1+e^{\beta(E_{n}-E_{m}+\mu)}}{E_{n}-E_{m}+\mu+i\frac{\pi}{\beta}(2r+1)},$ (30)
and hence
$\bar{G}_{\sigma}(\vec{k},i\frac{\pi}{\beta}|\mu)=\sum_{n,m}e^{\beta(\mu N_{n}-E_{n})}\frac{1+e^{\beta(E_{n}-E_{m}+\mu)}}{E_{n}-E_{m}+\mu+i\frac{\pi}{\beta}}|\langle n|C_{\sigma}(\vec{k})|m\rangle|^{2}.$ (31)
The reality of the eigenvalues $E_{n}$ shows that $\bar{G}$ has no poles in the strip Eq. (8) due to the cancellation of a possible pole at $E_{n}-E_{m}+\mu+i\frac{\pi}{\beta}=0$ by the vanishing of the numerator $1+e^{\beta(E_{n}-E_{m}+\mu)}$ at the same location. Noting also that $0\leq|\langle n|C|m\rangle|^{2}\leq 1$, $\bar{G}$ is therefore a holomorphic function of $\mu$ in the strip ${\cal S}$. From Eq. (26) we obtain the theorem.
4.1 Another proof of Eq. (2326)
We note a spectral representation for $\bar{A}$ (of the same form as Eq. (29))
$\bar{A}_{\sigma}(\vec{k},i\omega_{r})=\sum_{n,m}\frac{e^{-\beta\epsilon_{n}}+e^{-\beta\epsilon_{m}}}{\epsilon_{n}-\epsilon_{m}+i\frac{\pi}{\beta}(2r+1)}\langle n|a_{\sigma}(\vec{k})|m\rangle\langle m|C^{\dagger}_{\sigma}(\vec{k})|n\rangle.$ (32)
Using this relation and Eq. (29) we rearrange Eq. (21) with $r=0$ as
$Z(\mu)$ $=$ $\left\{i\frac{\pi}{\beta}+\mu-\varepsilon_{{\sigma}}(\vec{k})\right\}\sum_{nm}\frac{e^{-\beta\epsilon_{n}}+e^{-\beta\epsilon_{m}}}{\epsilon_{n}-\epsilon_{m}+i\frac{\pi}{\beta}}|\langle n|C_{\sigma}(\vec{k})|m\rangle|^{2}$ (33)
$-\sum_{n,m}\frac{e^{-\beta\epsilon_{n}}+e^{-\beta\epsilon_{m}}}{\epsilon_{n}-\epsilon_{m}+i\frac{\pi}{\beta}}\langle n|a_{\sigma}(\vec{k})|m\rangle\langle m|C^{\dagger}_{\sigma}(\vec{k})|n\rangle$
which can be simplified using
$\langle n|a_{\sigma}(\vec{k})|m\rangle$ $=$ $\langle n[C_{\sigma}(\vec{k}),{\cal H]}|m\rangle-(\varepsilon_{\sigma}(\vec{k})-\mu)\langle n|C_{\sigma}(\vec{k})|m\rangle,$ (34)
$=$ $(\epsilon_{m}-\epsilon_{n}+\mu-\varepsilon_{\sigma}(\vec{k}))\langle n|C_{\sigma}(\vec{k})|m\rangle.$
so that combining the two terms Eq. (3334), and simplifying gives
$Z(\mu)$ $=$ $\sum_{nm}\sum_{nm}(e^{-\beta\epsilon_{n}}+e^{-\beta\epsilon_{m}})|\langle n|C_{\sigma}(\vec{k})|m\rangle|^{2}$ (35)
$=$ $\mbox{Tr}e^{-\beta{\cal H}}\{C_{\sigma}(\vec{k}),C^{\dagger}_{\sigma}(\vec{k})\}$
$=$ $\mbox{Tr}e^{-\beta{\cal H}},$
i.e. we recover the exact definition of $Z$ Eq. (6). Summarizing, we see that Eq. (2326) consistently represents the vanishing of the partition function $Z(\mu)$ Eq. (6).
5 Some illustrative examples
5.1 The Non-interacting Fermi system
We note that the non-interacting Fermi model with a partition function
$Z_{non}(\mu)=\prod_{\vec{k}{\sigma}}\left(1+\frac{e^{\beta\mu}}{e^{\beta\varepsilon_{\sigma}(\vec{k})}}\right),$ (36)
provides a simple illustration of our ideas, since a vanishing condition of the non-interacting Fermion Green’s function
$G^{-1}_{(0){\sigma}}(\vec{k},i\frac{\pi}{\beta}|\mu)={i\frac{\pi}{\beta}+\mu-\varepsilon_{\sigma}(\vec{k})},$ (37)
gives a zero of $Z_{non}(\mu)$. The added imaginary part $i\frac{\pi}{\beta}$ to $\mu$ converts the Fermi factor into a Bose factor. It is therefore amusing to note that root finding is performed on the “Bosonized” partition function.
5.2 Exact Green’s function in the Atomic Limit
We study a less trivial example of a single site Hubbard model in the so called atomic limit, where electrons are localized on an atom. Here the partition function, its roots and the exact Greens function are all calculable explicitly, and show that the theorem applies well to the exact solution for $G$ and also enables us to check the perturbative solution- which is carried out to ${\cal O}(U^{2})$.
We consider the atomic limit of the Hubbard model with spin dependent energies $\varepsilon_{i{\sigma}}$
$H=\sum_{i{\sigma}}\varepsilon_{i{\sigma}}C^{\dagger}_{i{\sigma}}C_{i{\sigma}}+U\sum_{i}C^{\dagger}_{i\uparrow}C_{i\uparrow}C^{\dagger}_{i\downarrow}C_{i\downarrow},$ (38)
which is uncoupled in the sites “i” and therefore easily solvable. Let us first consider the partition function which is easily evaluated in terms of the fugacity $z$ and other parameters as
$Z(\mu)$ $=$ $\prod_{i=1,N_{s}}{\cal Z}_{i}(\mu),$
${\cal Z}_{i}(\mu)$ $=$ $1+z(e^{-\beta\varepsilon_{i\uparrow}}+e^{-\beta\varepsilon_{i\downarrow}})+z^{2}e^{-\beta(U+\varepsilon_{i\uparrow}+\varepsilon_{i\downarrow})}.$ (39)
We simplify the notation using $\varepsilon_{i{\sigma}}=\varepsilon_{i0}+{\sigma}\delta_{\varepsilon}$, so that
${\cal Z}_{i}(\mu)$ $=$ $1+2ze^{-\beta\varepsilon_{i0}}\cosh\beta\delta_{\varepsilon}+z^{2}e^{-\beta(U+2\varepsilon_{i0})},$ (40)
so that the two roots of ${\cal Z}_{i}(\mu)$ in the fugacity variable $z=e^{\beta\mu}$, which are denoted as $z_{i{\sigma}}$, can be calculated as
$z_{i{\sigma}}=-1\times e^{\beta(U+\varepsilon_{i0})}\left(\cosh\beta\delta_{\varepsilon}+{\sigma}\{\sinh^{2}(\beta\delta_{\varepsilon})+1-e^{-\beta U}\}^{\frac{1}{2}}\right).$ (41)
At $U=0$ the roots are $z_{i{\sigma}}=-e^{\beta\varepsilon_{i{\sigma}}}$ the free particle roots, while for non-zero $U$ the roots are either on the negative real $z$ line ($U$ positive) or off the $z$ real line ($U$ negative). We can also express this result in terms of virtual energies $\xi_{i{\sigma}}$ defined in Eq. (10) as
$\xi_{i{\sigma}}=\frac{U}{2}+\varepsilon_{i0}+\frac{1}{\beta}\log\left(\cosh\beta\delta_{\varepsilon}+{\sigma}\{\sinh^{2}(\beta\delta_{\varepsilon})+1-e^{-\beta U}\}^{\frac{1}{2}}\right)$ (42)
We now drop the site index $i$, and compute the average number of particles using $n_{\sigma}=-\frac{1}{\beta}\frac{\partial}{\partial\varepsilon_{\sigma}}\log{\cal Z}$
$n_{{\sigma}}(\mu)=\frac{1}{{\cal Z}(\mu)}\left(e^{\beta(\mu-\varepsilon_{\sigma})}+e^{\beta(2\mu-U-2\varepsilon_{0})}\right)$ (43)
Now we write down the exact atomic limit Greens function- known from Hubbard’s work[23]
$G_{{\sigma}}(i\omega_{n}|\mu)=\frac{1-n_{{\bar{\sigma}}}(\mu)}{i\omega_{n}+\mu-\varepsilon_{\sigma}}+\frac{n_{{\bar{\sigma}}}(\mu)}{i\omega_{n}+\mu-U-\varepsilon_{\sigma}},$ (44)
which can be rewritten in the form of a single term with a self energy
$G_{{\sigma}}(i\omega_{n}|\mu)$ $=$ $\frac{1}{i\omega_{n}+\mu-\varepsilon_{\sigma}-\Sigma_{\sigma}(i\omega_{n}|\mu)},$
$\Sigma_{\sigma}(i\omega_{n}|\mu)$ $=$ $Un_{{\bar{\sigma}}}+\frac{U^{2}n_{{\bar{\sigma}}}(1-n_{\bar{\sigma}})}{i\omega_{n}+\mu-\varepsilon_{\sigma}-U(1-n_{{\bar{\sigma}}})}$ (45)
We alternately calculate $n_{\sigma}$ from Eq. (44) using the familiar formula $n_{\sigma}=\frac{1}{\beta}\sum_{n}e^{i\omega_{n}0^{+}}G_{\sigma}(i\omega_{n})$. After some algebra one can verify that the result is the same as in Eq. (43).
Note also that the un-normalized Greens function Eq. (17) is given by
$\bar{G}_{\sigma}(i\omega_{n}|\mu)=\frac{1+e^{\beta(\mu-\varepsilon_{\sigma})}}{i\omega_{n}+\mu-\varepsilon_{\sigma}}+\frac{e^{\beta(\mu-\varepsilon_{\bar{\sigma}})}\left(1+e^{\beta(\mu-\varepsilon_{\sigma}-U)}\right)}{i\omega_{n}+\mu-\varepsilon_{\sigma}-U}.$ (46)
It is easily seen that $\bar{G}_{\sigma}(i\omega_{0})$ is a holomorphic function of $\mu$ in the strip ${\cal S}$ Eq. (8) due to the vanishing of the numerator at every location of the vanishing denominators.
5.2.1 Roots for small U: Non degenerate case $\delta_{\varepsilon}>0$
We find by expanding Eq. (42)
$\xi_{\uparrow}=\epsilon_{0}+\delta_{\varepsilon}+\frac{U}{2}{\coth\beta\delta_{\varepsilon}}-\beta U^{2}\frac{1}{8}\frac{\cosh\beta\delta_{\varepsilon}}{\sinh^{3}\beta\delta_{\varepsilon}}+{\cal O}(U^{3})$ (47)
We can find $\xi_{\downarrow}$ from the above by setting $\delta_{\varepsilon}\to-\delta_{\varepsilon}$.
5.2.2 Roots for small U: Degenerate case $\delta_{\varepsilon}=0$
Expanding the $\delta_{\varepsilon}=0$ limit of Eq. (42) we get
$\xi_{+}$ $=$ $\varepsilon_{0}-\sqrt{\frac{U}{\beta}}+\frac{\sqrt{\beta}}{12}U^{\frac{3}{2}}+{\cal O}(U^{\frac{5}{2}})$
$\xi_{-}$ $=$ $\varepsilon_{0}-\sqrt{\frac{U}{\beta}}-\frac{\sqrt{\beta}}{12}U^{\frac{3}{2}}+{\cal O}(U^{\frac{5}{2}})$ (48)
5.2.3 Perturbative Roots from $G$-The non-degenerate case
We next show how the small U roots given in Eq. (4748) can be obtained using the perturbative expansion of the Dyson self energy, as remarked below the theorem, i.e. the condition $\Psi_{\vec{k}{\sigma}}(\mu)=0$. Recall that
$G^{-1}_{\sigma}(i\frac{\pi}{\beta}|-i\frac{\pi}{\beta}+x)=x-\varepsilon_{\sigma}-\Sigma_{\sigma}(x)$ (49)
with
$\Sigma_{\sigma}(x)=Un_{{\bar{\sigma}}}+\frac{U^{2}n_{{\bar{\sigma}}}(1-n_{\bar{\sigma}})}{x-\varepsilon_{\sigma}-U(1-n_{{\bar{\sigma}}})}.$ (50)
For the purpose of this calculation we use the convenient variable $x$ which is related to the virtual energy through $\xi=x-\frac{U}{2}$. We next expand the exact occupation $n_{\sigma}$ appearing above, using Eq. (43), in a series in $U$. In terms of the Fermi function
$f_{z,{\sigma}}$ $=$ $\frac{z}{z+e^{\beta\varepsilon_{\sigma}}}$ (51)
we find
$n_{\sigma}$ $=$ $f_{\sigma}-\beta Uf_{\sigma}f_{\bar{\sigma}}(1-f_{\bar{\sigma}})+{\cal O}(U^{2})$ (52)
where ${\bar{\sigma}}=-{\sigma}$. We setup the root equation to ${\cal O}(U^{2})$ for $G_{\sigma}^{-1}=0$. Recall we are temporarily using $x=i\frac{\pi}{\beta}+\mu$, so that $z\to-e^{\beta x}$, and in terms of $x$
$f_{z,{\sigma}}|_{z\to-e^{\beta x}}$$\equiv$ $f_{\sigma}(x)=\frac{e^{\beta x}}{e^{\beta x}-e^{\beta\varepsilon_{\sigma}}}$
$f^{\prime}(x)$ $=$ $\beta f_{\sigma}(x)(1-f_{\sigma}(x))$ (53)
Expanding $n_{\sigma}$ using Eq. (52), and truncating to ${\cal O}(U^{2})$ or from diagrammatics directly, we get
$\Sigma_{\sigma}(x)=Uf_{{\bar{\sigma}}}-\beta U^{2}f_{\bar{\sigma}}(x)(1-f_{\bar{\sigma}}(x))f_{\sigma}(x)+\frac{U^{2}f_{{\bar{\sigma}}}(1-f_{\bar{\sigma}})}{x-\varepsilon_{\sigma}}+{\cal O}(U^{3})$ (54)
In Eq. (49) we substitute the above, and so the condition for a root is
$x-\varepsilon_{\sigma}-Uf_{\bar{\sigma}}(x)+\beta U^{2}f_{\sigma}(x)f_{\bar{\sigma}}(x)(1-f_{\bar{\sigma}}(x))-U^{2}\frac{f_{\bar{\sigma}}(x)(1-f_{\bar{\sigma}}(x))}{(x-\varepsilon_{\sigma})}=0$ (55)
and for therefore for ${\sigma}=\uparrow$
$x$ $=$ $\varepsilon_{\uparrow}+Uf_{\downarrow}(x)-\beta U^{2}f_{\uparrow}(x)f_{\downarrow}(x)(1-f_{\downarrow}(x))+U^{2}\frac{f_{\downarrow}(x)(1-f_{\downarrow}(x))}{x-\varepsilon_{\uparrow}}=0$ (56)
$=$ $\varepsilon_{\uparrow}+Uf_{\downarrow}(x)-U^{2}\beta f_{\downarrow}(x)(1-f_{\downarrow}(x))\left[f_{\uparrow}(x)-\frac{k_{B}T}{x-\varepsilon_{\uparrow}}\right]$
We can solve iteratively
$x=x_{0}+Ux_{1}+U^{2}x_{2}+\ldots$ (57)
so that
$x_{0}$ $=$ $\varepsilon_{\uparrow}+\eta=\varepsilon_{0}+\delta_{\varepsilon}+\eta$ (58)
$x_{1}$ $=$ $f_{\downarrow}(x_{0})=\frac{e^{\delta_{\varepsilon}}}{2\sinh\beta\delta_{\varepsilon}}$ (59)
where we set $\eta=0$ at the end. To get the full second order result, we need to expand
$f_{\downarrow}(x_{0}+Ux_{1})$ $=$ $f_{\downarrow}(x_{0})+Ux_{1}f^{\prime}_{\downarrow}(x_{0})+..$ (60)
$=$ $f_{\downarrow}(x_{0})+\beta Uf^{2}_{\downarrow}(x_{0})(1-f_{\downarrow}(x_{0}))+..$
Hence
$x_{2}=\beta f^{2}_{\downarrow}(x_{0})(1-f_{\downarrow}(x_{0}))-U^{2}\beta f_{\downarrow}(x_{0})(1-f_{\downarrow}(x_{0}))\left[f_{\uparrow}(x_{0}+\eta)-\frac{k_{B}T}{x_{0}-\varepsilon_{\uparrow}+\eta}\right]$ (61)
The term in square brackets on the right equals $\frac{1}{2}$ by taking the limit $\eta\to 0$. On simplification this leads to
$x_{2}=-\frac{\beta}{8}\frac{\cosh\beta\delta_{\varepsilon}}{\sinh^{3}(\beta\delta_{\varepsilon})}$ (62)
We therefore verify that Eq. (585962) combine to give the first three terms obtained directly Eq. (47).
5.2.4 Degenerate case
For the degenerate case $\varepsilon_{\sigma}=\varepsilon_{0}$ (i.e. $\delta_{\varepsilon}=0$), we show how the leading behaviour in Eq. (48) is obtained from the condition $\Psi_{\vec{k}{\sigma}}(\mu)=0$ from the theorem. We write the leading, i.e. ${\cal O}(U)$ terms of Eq. (55)
$x-\varepsilon_{0}=U\frac{1}{1-e^{\beta(\varepsilon_{0}-x)}}+\ldots,$ (63)
and expand the Bose factor for $x\sim\varepsilon_{0}$ and rearrange to obtain the leading solution
$x-\varepsilon_{0}=\pm\sqrt{\frac{U}{\beta}}+..$ (64)
To get the next correction we note that the right-hand side of Eq. (63) is expressible as
$\frac{U}{\beta(x-\varepsilon_{0})(1-\frac{1}{2}\beta(x-\varepsilon_{0})+\ldots)},$
so in the spirit of a perturbative expansion we substitute the leading result Eq. (64) in the denominator and expanding we get the next correction as
$x_{\pm}=\varepsilon_{0}\pm\sqrt{\frac{U}{\beta}}+\frac{U}{2}+{\cal O}(U^{3/2}),$ (65)
giving the first few terms in Eq. (48).
This expression shows that the doubly degenerate root of the U=0 system evolves into a pair of roots at non-zero U. The pair of roots are real for $U>0$ and complex-conjugates for $U<0$. The reader will notice a similarity between this result, and the results from the heuristic argument presented in Eq. (23) in the Introduction.
5.2.5 Satisfying the two sumrules Eq. (1415)
We note that the perturbative solutions Eq. (474865) satisfy the first sumrule Eq. (14), one checks that terms of orders other than ${\cal O}(U)$ cancel out. The solutions Eq. (474865) also satisfy the second sumrule Eq. (15), here terms of ${\cal O}(U^{n})$ with $n\neq 0$ cancel out. These sumrules are therefore of interest in numerical schemes, providing nontrivial constraints.
6 Virtual energies from self energy to ${\cal O}(U)$.
The self energy of the Hubbard model on $N_{s}$ sites to ${\cal O}(U)$ is readily obtained from standard perturbation theory [17] as
$\Sigma_{\sigma}(\vec{k},i\omega_{n})=\frac{U}{N_{s}}\sum_{\vec{p}}\frac{1}{e^{\beta(\varepsilon_{p{\bar{\sigma}}}-\mu)}+1},$ (66)
where ${\bar{\sigma}}=-{\sigma}$. Substituting into Eq. (28) we get the equation determining the virtual energy as
$\xi_{\vec{k}{\sigma}}$ $=$ $\varepsilon_{\vec{k}{\sigma}}-\frac{U}{2}+\frac{U}{N_{s}}\sum_{p}\frac{1}{1-e^{\beta(\varepsilon_{p{\bar{\sigma}}}-\xi_{k{\sigma}}-\frac{U}{2})}}+{\cal O}(U^{2}).$ (67)
Below we will present the numerical solutions of a simplified version
$\xi_{\vec{k}{\sigma}}$ $=$ $\varepsilon_{\vec{k}{\sigma}}+\frac{U}{2N_{s}}\sum_{p}\coth\frac{\beta}{2}(\xi_{k{\sigma}}-\varepsilon_{p{\bar{\sigma}}}),$ (68)
where we dropped a term of the ${\cal O}(U^{2})$ arising from expanding the $\coth$. The dropped term can be added back into other second order terms in the self energy. From the structure of Eq. (68) and since $\sum_{p}\varepsilon_{p{\bar{\sigma}}}=0$, we see that the sum-rule Eq. (14) is automatically satisfied.
[…]
In Fig. (2) we compare the virtual energies for a few typical values of positive $U$ found by solving Eq. (68) in the left panel, and by numerically finding the roots of the partition function in the right panel.
[…]
In Fig. (3) we compare the exact virtual energies and the approximate ones from Eq. (68) at four typical values of $U$. For smaller $U\leq 0.25$ the results are quite close. They suggest that it is reasonable to label the numerical solutions with the $k$ values read off from the approximate solutions. For the larger values of $U$ the agreement is poorer, although the extremities show a closer convergence.
Figure 3: Comparison between the exact virtual energies found from the partition function zeros (red dots) and the solutions of Eq. (68) (blue dots) at four different repulsive values of $U$ marked in the plots.
In Fig. (4) we display results for the attractive case, at a few negative values of $U$. All virtual energies are non-real in these calculations, requiring only a negative sign of $U$. The solutions of Eq. (68) capture the shrinking range of the real part of the virtual energies at the extremities in all cases. The quantitative agreement with the imaginary parts is not as good.
The formation of bound (Cooper) pairs in this case is similar to that of electron-electron binding posited in Dyson’s well known argument[19] about the instability of quantum electrodyamics with a flipped sign of $e^{2}$. The resulting non-convergence of perturbation theory in field theory, is based on the instability of the ground state due to pair formation, under the assumption of an infinite system. For the conventionally measured variables (energy, susceptibilities etc) related effects can only be seen in the limit of a large number of particles- i.e. in the thermodynamic limit. In the present case, we see that the distinct behaviour of the virtual states between the repulsive and attractive cases, namely the energies going from real values for repulsion to complex pairs for attraction. This striking change is visible even for small systems.
Figure 4: Attractive interactions All the attractive cases lead to complex virtual energies, as expected from arguments given in the text. Complex plots of the virtual energies with the real (imaginary) part on the x (y) axis. We compare the exact virtual energies found from the partition function zeros (red dots) and the solutions of Eq. (68) (blue dots) at four different attractive values of $U$ marked in the plots. We picture the degenerate roots by introducing a slight displacement in the real part. The formation of bound (Cooper type) pairs for $U<0$ shrink the bandwidth of the real part, these plots show such a tendency already at $U$=-$0.5$.
7 Concluding remarks
We have argued that the roots of the grand partition function of many-body models, such as the Hubbard model, can be found perturbatively in the interaction. Our method gives virtual energies $\xi_{\vec{k}{\sigma}}$, which are assigned (crystal) momentum and spin labels. The condition Eq. (28), can be used to systematically estimate the virtual energies $\xi_{\vec{k}{\sigma}}$, starting from the non-interacting values $\xi_{\vec{k}{\sigma}}|_{U\to 0}\to\varepsilon_{\sigma}(\vec{k})$. We have illustrated this idea with two examples (i) The single site Hubbard model, by solving for the thermodynamically derived virtual energies by perturbation theory to second order in U and compared with expansion in $U$ of the exact values that are readily found (ii) the Benzene molecule, i.e. a 6-site Hubbard ring by solving for the virtual energies in a 6-site Hubbard model by first order perturbation theory in $U$ of either sign and compared with the exact numerical results.
For repulsive U, the virtual energies are real in the lowest order calculations. Assuming that a substantial fraction of the virtual roots are real at low $T$ in certain energy ranges, their level density should control the compressibility, heat capacity and susceptibility for the corresponding densities. In this sense the methods developed here may be expected to connect with the Landau Fermi liquid and its many variants[24]. Towards this end it seems important to study the virtual energy distribution for larger systems, and to study its variation with the strength of the interaction $U$, and the possible signatures of a Mott-Hubbard gap. For the case of attractive interaction, the results presented here give non-real virtual energies, already to first order in U and for very small lattice sizes. Further applications of these ideas will be published in a forthcoming publication[25].
8 Acknowledgements
I thank Professor Elliott Lieb for helpful discussions in 2018 at Princeton. I also thank Muhammad Sedik for helpful comments.
Appendix A Zeros from the Luttinger-Ward formula
In this section we discuss another expression for the partition function of a many-body system found in the well-known work of Luttinger and Ward (LW)[20]. LW found a formally exact expression for the thermodynamic potential $\Omega$ (and therefore the partition function) from perturbation theory to all orders in the coupling. The LW result follows from a clever rearrangement of the thermal perturbation theory, whose convergence is implicitly assumed to be sufficiently “benign”. Their result for the grand potential $\Omega$
$\Omega=-\frac{1}{\beta}\log Z(\mu)$ (69)
can be written in terms of the partition function in the following way:
$Z(\mu)=Z_{A}(\mu)\times Z_{B}(\mu)\times Z_{C}(\mu)$ (70)
where
$\log Z_{A}(\mu)$ $=$ $\sum_{\vec{k}\,r{\sigma}}e^{i\omega_{r}0^{+}}\;\log[-G_{\sigma}^{-1}(\vec{k},i\omega_{r})]$ (71)
$\log Z_{B}(\mu)$ $=$ $\sum_{\vec{k}\,r{\sigma}}e^{i\omega_{r}0^{+}}\{G_{0}^{-1}(\vec{k},i\omega_{r})G_{\sigma}(\vec{k},i\omega_{r})-1\}$ (72)
and
$\log Z_{C}(\mu)=-\sum_{\nu\geq 1}\sum_{\vec{k}\,r{\sigma}}e^{i\omega_{r}0^{+}}\frac{1}{2\nu}\widetilde{\Sigma}^{(\nu)}_{\sigma}(\vec{k},i\omega_{r})\;G_{\sigma}(\vec{k},i\omega_{r})$ (73)
where $G_{0}$ is the non-interacting Greens function, $\widetilde{\Sigma}^{(\nu)}_{\sigma}(\vec{k},i\omega_{r})$ is the total $\nu$th order skeleton self-energy part, where the order of the diagram $\nu$ counts all explicit powers of the interaction, say U. In these expressions the term $e^{i\omega_{r}0^{+}}$ is required for convergence of the sum. Note that $Z_{B}$ and $Z_{C}$ vanish in the limit of free Fermions. Readers might recognize that $\log Z_{C}$ Eq. (73) is the Luttinger-Ward functional of $G$, with the property that its functional derivative gives the exact self energy, and thereby $\Omega$ is a stationary functional of $G_{\vec{k}{\sigma}}$.
From Eq. (71) we obtain the product form
$Z_{A}(\mu)$ $=$ $\prod_{\vec{k}{\sigma}\,r}e^{i\omega_{r}0^{+}}\;\{-G_{\sigma}^{-1}(\vec{k},i\omega_{r})\},$ (74)
$=$ $\prod_{\vec{k}{\sigma}}(-1)\Psi_{\vec{k}{\sigma}}(\mu)\times\left(\prod_{\vec{k}{\sigma}\,r\neq 0}e^{i\omega_{r}0^{+}}\;\{-G_{\sigma}^{-1}(\vec{k},i\omega_{r}|\mu)\}\right)$
using Eq. (25). Indeed the complete partition function $Z_{0}$ of the free Fermi theory is of this form with $G_{0}$ replacing $G$. It is seen from this expression that $Z_{A}$ (and hence $Z(\mu)$) contains the factor $\Xi(\mu)$ given in Eq. (27), and their quotient is expressible as a functional of $G$. The quotient is required to be a holomorphic function of $\mu$ in the strip ${\cal S}$, with no other zeros. This result not easy to prove directly, but must be true if all the zeros are given by the function $\Xi(\mu)$ given in Eq. (27).

### v3

@@
 Abstract
-We present a method for calculating the Yang-Lee partition function zeros of a translationally invariant model of lattice fermions, exemplified by the Hubbard model.
-The method rests on a theorem involving the single electron self-energy $\Sigma_{\sigma}(\vec{k},i\omega_{n})$ in the imaginary time Matsubara formulation.
+We present a new method for calculating the Yang-Lee partition function zeros of a translationally invariant model of lattice fermions, exemplified by the Hubbard model.
+The method rests on a theorem involving the usual single electron self-energy $\Sigma_{\sigma}(\vec{k},i\omega_{n}|\mu)$ with chemical potential $\mu$, in the imaginary time Matsubara formulation.
 The theorem maps the Yang-Lee zeros to a set of wavevector and spin labeled virtual energies $\xi_{\vec{k}{\sigma}}$.
-These, thermodynamically derived virtual energies, are solutions of equations involving the self-energy at corresponding $\vec{k}{\sigma}$’s.
+These, thermodynamically derived virtual energies, are solutions of the set of equations $\xi_{k{\sigma}}=\varepsilon_{\sigma}({\vec{k}})-\frac{1}{2}U+\Sigma_{\sigma}(\vec{k},i\frac{\pi}{\beta}|\xi_{k{\sigma}}+\frac{1}{2}U-i\frac{\pi}{\beta})=0$.
 Examples of the method in simplified situations are provided.
@@
 The currently existing methods unfortunately do not appear to be generalizable for antiferromagnets and for broader classes of systems, including the ones considered here.
-Our interest is in the study of quantum many-body systems, such as the Hubbard model.
-Locating the zeros and extracting meaningful information for these systems has proceeded at a sedate pace.
+Our interest is in the study of quantum many-body systems, such as the Hubbard model (see Eq.
+(5)).
+Locating the zeros and extracting meaningful information for these physically important systems has proceeded at a sedate pace.
 Here the major difficulty is the absence of systematic analytical methods.
@@
 However the full spectrum of zeros seems difficult to obtain from QMC.
+Studies of related problems employing tensor networks show promise in overcoming some of these limitations [13, 14].
 There have also been recent studies of partition function zeros for related quantum many-body models in a variety of contexts, these include decoherence studies[15], quantum quenches[16] and in quantum computation [17].
@@
 (5))- in any dimension.
-This method requires a knowledge of the self energy in the Matsubara imaginary time framework[16].
-Perturbative knowledge of the self-energy may be used to give approximate location of the zeros.
+This method requires a knowledge of the self-energy in the Matsubara imaginary time framework[18].
+While the self-energy is a highly non-trivial object to compute, a perturbative knowledge of the self-energy is available and may be used to give approximate location of the zeros.
 The method generalizes to nearby models in a straightforward way.
@@
 Here the role of lifting of degeneracy of the single site energy is also studied in detail.
-The exact self energy is expanded to ${\cal O}(U^{2})$, and to this order the perturbative results for $\xi$ are shown to agree with the exact ones.
+The exact self-energy is expanded to ${\cal O}(U^{2})$, and to this order the perturbative results for $\xi$ are shown to agree with the exact ones.
 This somewhat long section is mainly addressed to readers interested in details, and may be skipped by others.
-In Section(6) we use the first order (Hartree) approximation to the self energy and extract the virtual energies $\xi_{\vec{k}{\sigma}}$ to ${\cal O}(U)$ for both signs of the interaction for the 1-d Hubbard ring of length 6.
+In Section(6) we use the first order (Hartree) approximation to the self-energy and extract the virtual energies $\xi_{\vec{k}{\sigma}}$ to ${\cal O}(U)$ for both signs of the interaction for the 1-d Hubbard ring of length 6.
 This is familiar as the idealized Benzene molecule.
@@
 In Appendix(A) we recall the well known results of Luttinger and Ward[23] for the grand partition function.
-Their result expresses the result of an all-orders summation of perturbation theory.
+Their result neatly expresses the result of an all-orders summation of perturbation theory.
 We observe that the key function $\Xi(\mu)$ in Eq.
@@
 where $E_{n}$ and $N_{n}$ are the eigenvalues of $H$ and particle number for the state $n$ (detailed definitions are given after Eq.
-(29)), $Z_{N}$ is the canonical partition function for $N$ particles and $z$ is the fugacity
-$z(\mu)=e^{\beta\mu}.$ (7)
-$Z(\mu)$ is thus a polynomial in the fugacity $z$ of degree $2N_{s}$, and by the fundamental theorem of algebra, it has exactly $2N_{s}$ zeros in the complex $z$ plane- and hence expressible as $Z(\mu)=c_{0}\prod_{\alpha=1}^{2N_{s}}(z-z_{\alpha})$, where $z_{\alpha}$ are the Yang-Lee zeros of the given model.
+(30)), $Z_{N}$ is the canonical partition function for $N$ particles and $z$ is the fugacity $z(\mu)=e^{\beta\mu}.$ $Z(\mu)$ is thus a polynomial in the fugacity $z$ of degree $2N_{s}$, and by the fundamental theorem of algebra, it has exactly $2N_{s}$ zeros in the complex $z$ plane- and hence expressible as $Z(\mu)=c_{0}\prod_{\alpha=1}^{2N_{s}}(z-z_{\alpha})$, where $z_{\alpha}$ are the Yang-Lee zeros of the given model.
 We will find it more convenient to work in the complex $\mu$ plane rather than with the fugacity $z$.
@@
 We will write the partition function in the form
-$Z(\mu)=\prod_{\vec{k}{\sigma}}(1+\frac{e^{\beta\mu}}{e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{U}{2})}}).$ (9)
+$Z(\mu)=\prod_{\vec{k}{\sigma}}(1+\frac{e^{\beta\mu}}{e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{1}{2}U)}}).$ (8)
 With this convention- explained below- the roots of $Z$ occur at
-$\mu$$\to$ $-i\frac{\pi}{\beta}+\frac{U}{2}+\xi_{\vec{k}{\sigma}},\mbox{ or equivalently }$
-$z_{\alpha}$$\to$ $-e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{U}{2})}$ (10)
+$\mu$$\to$ $-i\frac{\pi}{\beta}+\frac{1}{2}U+\xi_{\vec{k}{\sigma}},\mbox{ or equivalently }$
+$z_{\alpha}$$\to$ $-e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{1}{2}U)}$ (9)
 where Eq.
@@
 We remark that while Eq.
-(9) has the same form as that of an ideal Fermi gas with energies $\xi_{\vec{k}{\sigma}}+\frac{U}{2}$, the latter are non-trivial complex functions of $T,U$.
+(8) has the same form as that of an ideal Fermi gas with energies $\xi_{\vec{k}{\sigma}}+\frac{1}{2}U$, the latter are non-trivial complex functions of $T,U$.
 For an interacting system $\xi_{\vec{k}{\sigma}}$ are not viewed as eigenvalues of any Hermitean Hamiltonian, and are allowed to be non-real.
@@
 (7), and the range of the imaginary part of $\xi$ in Eq.
-(11), are such that complex conjugate roots in the fugacity variable $z$ Eq.
-(7) lead to complex conjugate energies $\xi$.
+(10), are such that complex conjugate roots in the fugacity variable $z=e^{\beta\mu}$ lead to complex conjugate energies $\xi$.
 The partition function being a polynomial in $z$ with real coefficients, the roots in $z$ are either real or come in complex conjugate pairs.
@@
 (12) by summing over all $\{\vec{k}{\sigma}\}$.
-  • A second sumrule can be readily found by comparing the results of the single particle sector $N=1$, i.e the coefficient of $e^{\beta\mu}$ in Eq.
+  • A second sum-rule can be readily found by comparing the results of the single particle sector $N=1$, i.e the coefficient of $e^{\beta\mu}$ in Eq.
 (8).
@@
 We find the second sum-rule:
-    $\sum_{\vec{k}{\sigma}}e^{-\beta(\xi_{\vec{k}{\sigma}}+\frac{U}{2})}=\sum_{\vec{k}{\sigma}}e^{-\beta\varepsilon_{\vec{k}{\sigma}}}.$ (15)
+    $e^{-\beta\frac{1}{2}U}\sum_{\vec{k}{\sigma}}e^{-\beta\xi_{\vec{k}{\sigma}}}=\sum_{\vec{k}{\sigma}}e^{-\beta\varepsilon_{\vec{k}{\sigma}}}.$ (14)
+    We now define $U_{eff}$ as a function of the parameters $U,\beta$ and also the computed $\{\xi\}$. It is given by
+    $U_{eff}(U,\beta,\{\xi\})=\frac{2}{\beta}\log\left(\frac{\sum_{k{\sigma}}e^{-\beta\xi_{k{\sigma}}}}{\sum_{k{\sigma}}e^{-\beta\varepsilon_{k{\sigma}}}}\right).$ (15)
+    The sum-rule Eq. (14) may be cast in the form
+    $U_{eff}(U,\beta,\{\xi\})=U.$ (16)
+    Clearly if the $\xi$’s are exactly found, then the sum-rule is satisfied exactly.
+If $\xi$’s are only found approximately, then the quality of the approximation is reflected in the difference between $U_{eff}$ and $U$.
+This idea is illustrated below in Fig.
+(34).
   • Further sum-rules can be found by comparing other powers of $e^{\beta\mu}$, but they require exact knowledge of the canonical partition functions $Z_{2},Z_{3}\ldots$, and are therefore less tractable.
-3 The Greens function and Self energy in Matsubara frequencies
+3 The Greens function and Self-energy in Matsubara frequencies
 It is useful to consider the unnormalized thermal Greens function, signified by an overbar
@@
 $\left\{i\omega_{r}+\mu-\varepsilon_{\sigma}(k)\right\}\bar{G}_{\sigma}(\vec{k},i\omega_{r}|\mu)=Z(\mu)+\bar{\cal A}_{\sigma}(\vec{k},i\omega_{r}|\mu).$ (22)
-Defining the Dyson self energy $\Sigma$ as
+Defining the Dyson self-energy $\Sigma$ as
 $\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)=\frac{\bar{\cal A}_{\sigma}(\vec{k},i\omega_{r}|\mu)}{\bar{G}_{\sigma}(\vec{k},i\omega_{r}|\mu)}$ (23)
@@
 We also note that Eq.
-(302122) implies the periodicity in the complex $\mu$ plane for integer $s$
+(312223) implies the periodicity in the complex $\mu$ plane for integer ‘$s$
 $\bar{G}_{\sigma}(\vec{k},i\omega_{r}|\mu+is\frac{2\pi}{\beta})=\bar{G}_{\sigma}(\vec{k},i\omega_{r+s}|\mu),$ (25)
@@
   • In principle the self-energy $\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)$ can be calculated to arbitrary orders in the interaction, using the standard rules of the Feynman-Dyson-Matsubara perturbation theory in the imaginary time[18, 19, 23].
-Given an approximation for $\Sigma$ to a certain order in the coupling, the theorem can be used to find the corresponding (unlabeled) Yang-Lee zero of the partition function, or more precisely, the momentum-labeled virtual energy $\xi_{\vec{k}{\sigma}}$.
+Given an approximation for $\Sigma$ to a certain order in the coupling, the theorem can be used to find the corresponding momentum-labeled virtual energy $\xi_{\vec{k}{\sigma}}$, and from their total set, the (unlabeled) Yang-Lee zeros of the partition function.
   • If the non-interacting virtual energies (i.e.
@@
   • Using Eq. (9) and Eq. (26) we summarize the final equation determining the virtual energy $\xi_{\vec{k}{\sigma}}$ as
-    $\xi_{\vec{k}{\sigma}}=\varepsilon_{\vec{k}{\sigma}}-\frac{U}{2}+\Sigma_{{\sigma}}(\vec{k},i\frac{\pi}{\beta}|\xi_{k{\sigma}}+\frac{U}{2}-i\frac{\pi}{\beta}).$ (28)
+    $\xi_{\vec{k}{\sigma}}=\varepsilon_{\vec{k}{\sigma}}-\frac{1}{2}U+\Sigma_{{\sigma}}(\vec{k},i\frac{\pi}{\beta}|\xi_{k{\sigma}}+\frac{1}{2}U-i\frac{\pi}{\beta}).$ (29)
 Proof: To prove the theorem, a glance at Eq.
@@
 (9) as
-$\xi_{i{\sigma}}=\frac{U}{2}+\varepsilon_{i0}+\frac{1}{\beta}\log\left(\cosh\beta\delta_{\varepsilon}+{\sigma}\{\sinh^{2}(\beta\delta_{\varepsilon})+1-e^{-\beta U}\}^{\frac{1}{2}}\right)$ (42)
+$\xi_{i{\sigma}}=\frac{1}{2}U+\varepsilon_{i0}+\frac{1}{\beta}\log\left(\cosh\beta\delta_{\varepsilon}+{\sigma}\{\sinh^{2}(\beta\delta_{\varepsilon})+1-e^{-\beta U}\}^{\frac{1}{2}}\right)$ (43)
 We now drop the site index $i$, and compute the average number of particles using $n_{\sigma}=-\frac{1}{\beta}\frac{\partial}{\partial\varepsilon_{\sigma}}\log{\cal Z}$
@@
 $G_{{\sigma}}(i\omega_{n}|\mu)=\frac{1-n_{{\bar{\sigma}}}(\mu)}{i\omega_{n}+\mu-\varepsilon_{\sigma}}+\frac{n_{{\bar{\sigma}}}(\mu)}{i\omega_{n}+\mu-U-\varepsilon_{\sigma}},$ (45)
-which can be rewritten in the form of a single term with a self energy
+which can be rewritten in the form of a single term with a self-energy
 $G_{{\sigma}}(i\omega_{n}|\mu)$ $=$ $\frac{1}{i\omega_{n}+\mu-\varepsilon_{\sigma}-\Sigma_{\sigma}(i\omega_{n}|\mu)},$
@@
 It is easily seen that $\bar{G}_{\sigma}(i\omega_{0})$ is a holomorphic function of $\mu$ in the strip ${\cal S}$ Eq.
-(8) due to the vanishing of the numerator at every location of the vanishing denominators.
+(7), due to the vanishing of the numerator at every location of the vanishing denominators.
 5.2.1 Roots for small U: Non degenerate case $\delta_{\varepsilon}>0$
 We find by expanding Eq. (43)
-$\xi_{\uparrow}=\epsilon_{0}+\delta_{\varepsilon}+\frac{U}{2}{\coth\beta\delta_{\varepsilon}}-\beta U^{2}\frac{1}{8}\frac{\cosh\beta\delta_{\varepsilon}}{\sinh^{3}\beta\delta_{\varepsilon}}+{\cal O}(U^{3})$ (47)
+$\xi_{\uparrow}=\epsilon_{0}+\delta_{\varepsilon}+\frac{1}{2}U{\coth\beta\delta_{\varepsilon}}-\beta U^{2}\frac{1}{8}\frac{\cosh\beta\delta_{\varepsilon}}{\sinh^{3}\beta\delta_{\varepsilon}}+{\cal O}(U^{3})$ (48)
 We can find $\xi_{\downarrow}$ from the above by setting $\delta_{\varepsilon}\to-\delta_{\varepsilon}$.
@@
 We next show how the small U roots given in Eq.
-(4748) can be obtained using the perturbative expansion of the Dyson self energy, as remarked below the theorem, i.e.
+(4849) can be obtained using the perturbative expansion of the Dyson self-energy, as remarked below the theorem, i.e.
 the condition $\Psi_{\vec{k}{\sigma}}(\mu)=0$.
@@
 $\Sigma_{\sigma}(x)=Un_{{\bar{\sigma}}}+\frac{U^{2}n_{{\bar{\sigma}}}(1-n_{\bar{\sigma}})}{x-\varepsilon_{\sigma}-U(1-n_{{\bar{\sigma}}})}.$ (51)
-For the purpose of this calculation we use the convenient variable $x$ which is related to the virtual energy through $\xi=x-\frac{U}{2}$.
+For the purpose of this calculation we use the convenient variable $x$ which is related to the virtual energy through $\xi=x-\frac{1}{2}U$.
 We next expand the exact occupation $n_{\sigma}$ appearing above, using Eq.
@@
 so in the spirit of a perturbative expansion we substitute the leading result Eq. (65) in the denominator and expanding we get the next correction as
-$x_{\pm}=\varepsilon_{0}\pm\sqrt{\frac{U}{\beta}}+\frac{U}{2}+{\cal O}(U^{3/2}),$ (65)
+$x_{\pm}=\varepsilon_{0}\pm\sqrt{\frac{U}{\beta}}+\frac{1}{2}U+{\cal O}(U^{3/2}),$ (66)
 giving the first few terms in Eq. (49).
@@
 These sumrules are therefore of interest in numerical schemes, providing nontrivial constraints.
-6 Virtual energies from self energy to ${\cal O}(U)$.
-The self energy of the Hubbard model on $N_{s}$ sites to ${\cal O}(U)$ is readily obtained from standard perturbation theory [17] as
+6 Virtual energies from self-energy to ${\cal O}(U)$.
+The self-energy of the Hubbard model on $N_{s}$ sites to ${\cal O}(U)$ is readily obtained from standard perturbation theory [19] as
 $\Sigma_{\sigma}(\vec{k},i\omega_{n})=\frac{U}{N_{s}}\sum_{\vec{p}}\frac{1}{e^{\beta(\varepsilon_{p{\bar{\sigma}}}-\mu)}+1},$ (67)
 where ${\bar{\sigma}}=-{\sigma}$. Substituting into Eq. (29) we get the equation determining the virtual energy as
-$\xi_{\vec{k}{\sigma}}$ $=$ $\varepsilon_{\vec{k}{\sigma}}-\frac{U}{2}+\frac{U}{N_{s}}\sum_{p}\frac{1}{1-e^{\beta(\varepsilon_{p{\bar{\sigma}}}-\xi_{k{\sigma}}-\frac{U}{2})}}+{\cal O}(U^{2}).$ (67)
+$\xi_{\vec{k}{\sigma}}$ $=$ $\varepsilon_{\vec{k}{\sigma}}-\frac{1}{2}U+\frac{U}{N_{s}}\sum_{p}\frac{1}{1-e^{\beta(\varepsilon_{p{\bar{\sigma}}}-\xi_{k{\sigma}}-\frac{1}{2}U)}}+{\cal O}(U^{2}).$ (68)
 Below we will present the numerical solutions of a simplified version
@@
 where we dropped a term of the ${\cal O}(U^{2})$ arising from expanding the $\coth$.
-The dropped term can be added back into other second order terms in the self energy.
+The dropped term is added back to other second order terms in the self-energy, and the results of such a calculation to ${\cal O}(U^{2})$ will be reported later [20].
 From the structure of Eq.
@@
 For the larger values of $U$ the agreement is poorer, although the extremities show a closer convergence.
-Figure 3: Comparison between the exact virtual energies found from the partition function zeros (red dots) and the solutions of Eq.
+Figure 3: Comparison between the exact virtual energies found from the partition function zeros (red dots) and the solutions of the first order equations Eq.
 (69) (blue dots) at four different repulsive values of $U$ marked in the plots.
+The $x$ axis label index “$\nu$” is as defined in Fig.
+(2) for the Hartree solution and also the sorting index for the energies found from ED.
+For the displayed $U$ and $\beta$, the solutions of Eq.
+(69) can be used to compute $U_{eff}$ (see Eq.
+(1516)).
+The pairs $(U,U_{eff})$ are $(0.25,0.149),(0.5,0.289),(1.0,0.55),(1.5,0.804)$, indicating that the first order approximation worsens with increasing $U$.
 In Fig.
@@
 The formation of bound (Cooper) pairs in this case is similar to that of electron-electron binding posited in Dyson’s well known argument[22] about the instability of quantum electrodyamics with a flipped sign of $e^{2}$.
-The resulting non-convergence of perturbation theory in field theory, is based on the instability of the ground state due to pair formation, under the assumption of an infinite system.
+The resulting non-convergence of perturbation theory in field theory, is based on the instability of the ground state due to pair formation, under the standard assumption of an infinite system.
 For the conventionally measured variables (energy, susceptibilities etc) related effects can only be seen in the limit of a large number of particles- i.e.
@@
 This striking change is visible even for small systems.
-Figure 4: Attractive interactions All the attractive cases lead to complex virtual energies, as expected from arguments given in the text.
-Complex plots of the virtual energies with the real (imaginary) part on the x (y) axis.
-We compare the exact virtual energies found from the partition function zeros (red dots) and the solutions of Eq.
-(68) (blue dots) at four different attractive values of $U$ marked in the plots.
-We picture the degenerate roots by introducing a slight displacement in the real part.
-The formation of bound (Cooper type) pairs for $U<0$ shrink the bandwidth of the real part, these plots show such a tendency already at $U$=-$0.5$.
+[…]
 7 Concluding remarks
@@
 In this sense the methods developed here may be expected to connect with the Landau Fermi liquid and its many variants[27].
-Towards this end it seems important to study the virtual energy distribution for larger systems, and to study its variation with the strength of the interaction $U$, and the possible signatures of a Mott-Hubbard gap.
+Towards this end it seems important to study the virtual energy distribution for larger systems, and to study its variation with the strength of the interaction $U$.
 For the case of attractive interaction, the results presented here give non-real virtual energies, already to first order in U and for very small lattice sizes.
@@
 In this section we discuss another expression for the partition function of a many-body system found in the well-known work of Luttinger and Ward (LW)[23].
-LW found a formally exact expression for the thermodynamic potential $\Omega$ (and therefore the partition function) from perturbation theory to all orders in the coupling.
+LW found a formally exact expression for the thermodynamic potential $\Omega$ (and therefore the partition function) in terms of the Greens function $G_{\sigma}(\vec{k},i\omega_{r})$ , from perturbation theory to all orders in the coupling.
 The LW result follows from a clever rearrangement of the thermal perturbation theory, whose convergence is implicitly assumed to be sufficiently “benign”.
@@
 $\log Z_{C}(\mu)=-\sum_{\nu\geq 1}\sum_{\vec{k}\,r{\sigma}}e^{i\omega_{r}0^{+}}\frac{1}{2\nu}\widetilde{\Sigma}^{(\nu)}_{\sigma}(\vec{k},i\omega_{r})\;G_{\sigma}(\vec{k},i\omega_{r})$ (74)
-where $G_{0}$ is the non-interacting Greens function, $\widetilde{\Sigma}^{(\nu)}_{\sigma}(\vec{k},i\omega_{r})$ is the total $\nu$th order skeleton self-energy part, where the order of the diagram $\nu$ counts all explicit powers of the interaction, say U.
+where $G_{0}$ is the non-interacting Greens function, $\widetilde{\Sigma}^{(\nu)}_{\sigma}(\vec{k},i\omega_{r})$ is the total $\nu$th order skeleton self-energy part, where the order of the diagram $\nu$ counts all explicit powers of the interaction, in this case U.
 In these expressions the term $e^{i\omega_{r}0^{+}}$ is required for convergence of the sum.
@@
 Readers might recognize that $\log Z_{C}$ Eq.
-(73) is the Luttinger-Ward functional of $G$, with the property that its functional derivative gives the exact self energy, and thereby $\Omega$ is a stationary functional of $G_{\vec{k}{\sigma}}$.
+(74) is the Luttinger-Ward functional of $G$, with the property that its functional derivative gives the exact self-energy, and thereby $\Omega$ is a stationary functional of $G_{\vec{k}{\sigma}}$.
 From Eq. (72) we obtain the product form